How many milliliters of a 6.00 M HCl solution are required to react with 21.0 g magnesium?
1
Expert's answer
2020-12-03T13:49:00-0500

n(Mg) = m/M = 21g / 24 g/mol = 0.875 mol


Mg + 2HCl = MgCl2 + H2


n(HCl)=2n(Mg)=0.875*2=1.75 mol

V=n/c=1.75 mol / 6 mol/L = 0.292 L = 292 mL


V(HCl solution) = 292 mL